A nurse is preparing to administer erythromycin estolate (Ilosone) 30 mg/kg PO divided in equal doses every 6 hr to a child who weighs 28lb. Available is erythromycin estolate suspension 125 mg/5 mL. How many mL should the nurse administer per dose? round to nearest tenth if applicable
Answer #1

The child's weight = 28 lbs = 12.7 kg

The dose for the day = 30mg/kg

The dose per day for the child = 12.7 x 30 = 381 mg

The daily dose has to be divided equally every 6 hourly, so every 6 hourly dose is

   = 381/4 = 95.25 mg

The dose in hand = 125 mg/5ml

How much ml has to be administered per dose =    dose prescribed x Amount in hand dose in hand

   = 95.25 x5 /125 = 3.81ml 0r 3.8 ml